You are here: Foswiki>PartnerObserving Web>ItalyQueue2012>ObsInaf20121206 (28 Feb 2013, EleanoraSani)Edit Attach

INAF LBC+MODS+LUCI queue observing, 2012 December 5-6

Observer: E. Sani, F. Cusano, M. Fumana
Telescope Operator: G. Becchetti
Support Astronomer: M. Edwars
Instrument: LUCI1

Summary

At the twilight scattered cirrus. Wind speed is 10m/s
At the beginning of the observations the average seeing is 0.6".
According to the S.A. infos, the rotator PA for the mos acquisition (for masks without reference boxes) could be not enough accurate.
During the night cirrus last but with average seeing 0.5"-0.6"

Data Summary

Project_RUN Objects Filter ExpTime Notes
Cresci_28A DSF2237a K, 210_zJHK 1h15m  
Pentericci_02 1UDS HKspec, 200_H+K 2h Tell+flat+arc
Andreon_01 JKCS041 HKspec, 200_H+K 1h Tell+flat+arc
Hunt_08 GRB020127 HKspec, 200_H+K 1h Tell+flat+arc
LaFranca_07 NGC4138 zJspec, 200_H+K 45m Tell. COMPLETED
LaFranca_11 2MSXJ11+19 zJspec, 200_H+K 45m Tell. COMPLETED
LaFranca_11 2MSXJ11+19 K, 210_ZJHK 40m Tell. COMPLETED

10.15 We perform dark calibrations
  • ./DoDarks luci.20121206.0001-0155
  • DoDarks starts with dark_int.txt, then dark_nor.txt

11:56 Dark script interupted. We need to open dome.

12.16 Dome Open

01.10 executeLuciscript.sh tell2_acq_campoA.txt luci.20121206.0156-0158

01.24 executeLuciscript.sh tell_int_campoA_K.txt luci.20121206.0159-0166

01:27 executeLuciscript.sh acq_campoA.txt =luci.20121206.0167-0173

01:53 executeLuciscript.sh sci_campoA_K_1.txt luci.20121206.0174-185
  • mean seeing is 0.6"
  • guide problem: guider timed out. Image luci.20121206.0185.fits can be affected by NO guiding
  • restarting gcs

03:00 we need to acquire again the scientific target problem with guiding we need to call John Hill to fix the problem

03:09 executeLuciscript.sh acq_campoA.txt luci.20121206.0186-0193

03:31 executeLuciscript.sh sci_campoA_K_1.txt luci.20121206.0194-096

03.51 Problem the LUCI software exits the scrip. NOT clear if it is a LUCI problem or a Guider problem. The error message on the GCS is "couldn't pause guide loop"
  • We verify that the last image luci.20121206.01946 is NOT affected by the problem, that appeared just while the telescope was performing the following offset
  • Error message from the IIFGUI Thu Dec 6 03:49:51.517 2012 Offset alert for left LUCIFER: guide loop didn't pause within timeout. Thu Dec 6 03:49:54.582 2012 Offset command failed for left LUCIFER

04.22 Cresci_28A is setting. We slew to Pentericci_02
  • executeLuciscript.sh pente1.acq luci.20121206.0197-0201
  • during the acquisition the seeing is 0.5"

04.47 executeLuciscript.sh pente1_HK_1.spec luci.20121206.0202-0213
  • Medium cirrus

05.54 executeLuciscript.sh pente1_HK_2.spec luci.20121206.0214-0223

06:52 executeLuciscript.sh andreon_mos.acq luci.20121206.0224-0226

07:07 executeLuciscript.sh andreon_HK_1.spec luci.20121206.0227-0238

08:30 Slew to Hunt_08

08:32 executeLuciscript.sh GRB020127acquisition_tel.script luci.20121206.0239-0240

08:38 executeLuciscript.sh GRB020127science_HKspec_tel.script luci.20121206.0241-0249

08:46 executeLuciscript.sh GRB020127acquisition.script luci.20121206.0250-0253

08:57 executeLuciscript.sh GRB020127science_HKspec.script luci.20121206.0254-265
  • Script interrupt after 1 offset (12 frame acquired)

10:02 Slew to La Franca
  • Wrong camera in script configuration (N3.5 instead of N1.8).

10:07 Slew to La Franca telluric
  • executeLuciscript.sh STDA7_1_J.acq luci.20121206.0266-0268

10:15 executeLuciscript.sh STDA7_1_200_zJspec.spec luci.20121206.0269-0270

10:18 Slew to La Franca target
  • executeLuciscript.sh OB_LaFranca_A7_acq.txt luci.20121206.0271-0274

10:29 executeLuciscript.sh OB_LaFranca_A7_spe.txt luci.20121206.0275-0282

11:22 Slew to La Franca 11
  • executeLuciscript.sh STDA11_1_J.acq luci.20121206.0283-0286

11.31 executeLuciscript.sh STDA11_1_200_zJspec.spec luci.20121206.0287-0288

11.34 executeLuciscript.sh STDA11_1_210_K.spec luci.20121206.0289-0290

11.38 executeLuciscript.sh A11_1_J.acq luci.20121206.291-0295
  • average seeing is 1"

11.49 executeLuciscript.sh A11_1_200_zJspec.spec luci.20121206.0296-302 *average seeing 0.6"

12.41 A11_1_210_K.spec luci.20121206.0303-0310 *spectrum integrated during the twilight but the background count remain stable for frames luci.20121206.0303-0308
  • background increase from about 16900 counts to 18500 in luci.20121206.0309
  • the last image is realized without guiding and only tracking luci.20121206.0310

01.35 The T.O. close the dome

01.45 We perform flats and arcs for the Sozzetti_22 program
  • executeLuciscript.sh spec_cal_lamp_22.text luci.20121206.0311-0320
  • this first arcs are wrong, we do not know why and acquire them by hand luci.20121206.0321-0330 we double the exposure time of the last five lamp acquisitions

-- VincenzoTesta - 05 Dec 2012
Topic revision: r12 - 28 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback